OpenStudy (angela210793):

Help me please...i want to sleep....it's 11:20pm in here....i've been looking 4 a solution for : b/a=1/3,a/c=2/3,a,b,c>0,a+b+c=34...find a=? and i still haven't found the right 1 (i guess)...

OpenStudy (amistre64):

you must be across the atlantic from me

OpenStudy (angela210793):

yea...told u I'm in Albania(in north of Greece) Europe...:)

OpenStudy (anonymous):

This is a tricky one. Let me try.

OpenStudy (angela210793):

sure...take ur time...i got 3*68/17...O.o

OpenStudy (amistre64):

b/a=1/3 a/c=2/3 ,a,b,c>0, a+b+c=34 ; /a maybe> 1/3 + c/a = 34/a ; a/c = 2/3..so flip it 1/3 + 3/2 = 34/a 2+9 ---- = 34/a 6 11 = 204/a a = 204/11 maybe :)

OpenStudy (anonymous):

I got 204/11 also.

OpenStudy (angela210793):

u did? O.o (maybe it's my fault cause I'm half asleep) but i didn't get it...

OpenStudy (angela210793):

let me see again

OpenStudy (amistre64):

b/a = 1/3 right? a/c = 2/3 ; so c/a = 3/2 a/a + b/a + c/a = 34/a opps; left out the 1 lol

OpenStudy (amistre64):

1+(1/3) + (3/2) = 34/a

OpenStudy (anonymous):

correct

OpenStudy (angela210793):

Ahhaaaa....now i get it....Thanks a looooooooooooooooooooot :):):)

OpenStudy (anonymous):

:)